OpenStudy (anonymous):

solve the system. {-5x - y = 21 3x - 4y +5z = -8 -3x -z = 12

OpenStudy (texaschic101):

are there any z's in the 1st equations or any y's in the 3rd equation ?

OpenStudy (texaschic101):

@nsschwab23 .....???

OpenStudy (anonymous):

no there is not

OpenStudy (anonymous):

y=-5x-21 ---> y=-1 3x+20x+84-15x-60=-8 ---> x=-4 z=-3x-12 -----> z=0

OpenStudy (texaschic101):

-5x - y = 21 --> -y = 5x + 21 --> y = -5x - 21 -3x - z = 12 --> -z = 3x + 12 --> z = - 3x - 12 now sub these into 2nd equation 3x - 4y + 5z = - 8 3x - 4(-5x - 21) + 5(-3x - 12) = - 8 3x + 20x + 84 - 15x - 60 = - 8 8x + 24 = - 8 8x = -8 - 24 8x = - 32 x = - 32/8 = - 4 now sub -4 in for x y = - 5x - 21 y = - 5(-4) - 21 y = 20 - 21 y = -1 go ahead and sub -4 into the 3rd equation z = -3x - 12 z = -3(-4) - 12 z = 12 - 12 z = 0 check... 3x - 4y + 5z = -8 3(-4) - 4(-1) + 5(0) = - 8 -12 + 4 + 0 = - 8 -8 = -8 (correct) ANSWER : x = - 4, y = - 1, z = 0

OpenStudy (anonymous):

thank you for the help

OpenStudy (texaschic101):

your welcome....glad to help :)